The length of a rectangular pipe is five

times longer than its width. Find the dimensions that give a​ cross-sectional area of 245

sq in

Answer by stanbon(75887)

The length of a rectangular pipe is five
times longer than its width. Find the dimensions that give a​ cross-sectional area of 245
-----
width:: w
length:: 5w
------
Equation:
w*5w = 245
5w^2 = 245
w^2 = 49
width = 7
length = 5*7 = 35
--------------
Cheers,
Stan H.
-------------